### v3

@@ 2.2 Binary relations
A binary relation $R^{\prime}$ on $A^{\prime}\subseteq A$ is a linear order extension of an asymmetric binary relation $R$ on $A^{\prime}$ if $R^{\prime}$ is a linear order on $A^{\prime}$ and $R\subseteq R^{\prime}$. Since $R$ is asymmetric, a linear order extension of $R$ is well-defined. Let $\mathcal{E}\left(R\right)$ be the set of all linear order extensions of $R$ on $A^{\prime}$. The following fact is also known.
@@
If $R$ on $A^{\prime}\subseteq A$ is acyclic, then $\mathcal{E}\left(R\right)\neq\emptyset$. [^5] If $A$ is an inifinite set, then acyclicity is not sufficient for the existence of linear order extenstions. In this case, Suzumura-consistency is a necessary and sufficient condition for the existence of linear order extenstions. See, Theorem 2.6 of Bossert and Suzumura (2010), which is a generalization of Szplrajn’s theorem. However, if $A$ is a finite set, then acyclicity is sufficient.
Note that since an acyclic binary relation is asymmetric, If $R$ on $A^{\prime}\subseteq A$ is acyclic, then a linear order extension $R$ is well-defined.
A linear order extension of an acyclic $R$ on $A^{\prime}$ is constructed as follows. First, choose one $a_{1}\in A^{\prime}$ such that $\left(a,a_{1}\right)\notin R$ for all $a\in A^{\prime}$. A result introduced by Bossert and Suzumura (2010, Theorem 2.4) ensures the existence of such $a_{1}$ with acyclic $R$. Similarly, for $t=2,\cdots,\left|A^{\prime}\right|$, we can let $a_{t}\in A^{\prime}\setminus\left\{a_{1},\cdots,a_{t-1}\right\}$ such that $\left(a,a_{t}\right)\notin R$ for all $a\in A^{\prime}\setminus\left\{a_{1},\cdots,a_{t-1}\right\}$. Then, $R:a_{1},a_{2},\cdots,a_{\left|A^{\prime}\right|}$ is a linear order extension of an acyclic $R$ on $A^{\prime}$. See also Kitahara and Okumura (2023) for this construction.
@@ 2.4 Graphs
We consider two cycles $C$ and $C^{\prime}$ of $G,$ where the nodes included in $C$ and $C^{\prime}$ are represented by $A$ and $A^{\prime}$, respectively. We say that $C$ is larger than $C^{\prime}$ if $A^{\prime}\subsetneq A$. A cycle $C$ of $G$ is said to be a maximal cycle of $G$ if there is no cycle $C^{\prime}$ of $G$ that is larger than $C$.
@@ 3 Results
$=$ $\{a_{2}a_{4},a_{2}a_{1},a_{2}a_{3},a_{4}a_{1},a_{4}a_{3},$
$a_{1}a_{3},a_{5}a_{6},a_{5}a_{4},a_{6}a_{4},a_{7}a_{6}\},$

### v4

@@ 2.2 Binary relations
A binary relation $R^{\prime}$ on $A^{\prime}\subseteq A$ is a linear order extension of an asymmetric binary relation $R$ on $A^{\prime}$ if $R^{\prime}$ is a linear order on $A^{\prime}$ and $R\subseteq R^{\prime}$. Since $R$ is asymmetric, a linear order extension of $R$ is well-defined. Let $\mathcal{E}\left(R\right)$ be the set of all linear order extensions of $R$ on $A^{\prime}$. The following fact is also known. See, for example, Duggan (1999) and Bossert and Suzumura (2010).
@@
If $R$ on $A^{\prime}\subseteq A$ is acyclic, then $\mathcal{E}\left(R\right)\neq\emptyset$.
@@ 2.4 Graphs
We consider two cycles $C^{\prime}$ and $C^{\prime\prime}$ of $G,$ where the nodes included in $C^{\prime}$ and $C^{\prime\prime}$ are represented by $A^{\prime}$ and $A^{\prime\prime}$, respectively. We say that $C^{\prime}$ is larger than $C^{\prime\prime}$ if $A^{\prime\prime}\subsetneq A^{\prime}$. A cycle $C$ of $G$ is said to be a maximal cycle of $G$ if there is no cycle $C^{\prime}$ of $G$ that is larger than $C$.
@@ 3 Results
$=$ $\{\left(a_{2},a_{4}\right),\left(a_{2},a_{1}\right),\left(a_{2},a_{3}\right),\left(a_{4},a_{1}\right),\left(a_{4},a_{3}\right),$
$\left(a_{1},a_{3}\right),\left(a_{5},a_{6}\right),\left(a_{5},a_{4}\right),\left(a_{6},a_{4}\right),\left(a_{7},a_{6}\right)\},$